x^6-2x^4-x^2+2=0 Am I supposed to factor by grouping here?
x^4(x^2-2) -1(x^2+2)
(x^4-1) (x^2-2) (x^2+2) Did I factor these correctly?
What would be my next step? Or did I miss a step?
I am totally lost with this problem. I am also having problems factoring terms that have higher than one as it's coefficient. Also, numbers with powers due to the zero product property being used.

Answer by richwmiller(17219)

x^6-2x^4-x^2+2=0
(x-1) (x+1) (x^2-2) (x^2+1) = 0